OpenStudy (anonymous):

limit x to inf of (1-(2/x))^(3x)

OpenStudy (anonymous):

\[\LARGE \lim_{x\to \infty} \;\left(1-\frac2x\right)^{3x}\] Do you mean this??

OpenStudy (anonymous):

yes thats what i mean, sorry for the way i typed it

OpenStudy (anonymous):

Never mind... ;) \[\LARGE \lim_{x\to \infty} \;\left(1-\frac{1}{\cfrac2x}\right)^{3x}= \lim_{x\to \infty} \;\left[\left(1-\frac{1}{\cfrac2x}\right)^{\frac2x\cdot \frac x2}\right]^{3x}\] \[\Large \lim_{x\to \infty} \;\left[\left(1-\frac{1}{\cfrac2x}\right)^{\frac2x}\right]^{\frac{3x^2}{2}}\] \[\Huge \lim_{x\to \infty}\;\; e^{\frac{3x^2}{2}}\] can you continue ??

OpenStudy (anonymous):

ahhh.. I made a mistake at principal... it's not 1/(2x) it is 1/(x/2) just replace that.. and follow the steps !

OpenStudy (anonymous):

i cant sorry. you mind finishing it for me?

OpenStudy (anonymous):

It's hard with latex... it takes time.. I have to go to school, after 30 minutes ! ...

OpenStudy (anonymous):

ask your teacher, he'll guide you ! Have nice day !

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):